Write maximum range in projectile motion.


1
Expert's answer
2021-02-13T12:18:09-0500

The maximum range in projectile motion can be written as follows:


"R=\dfrac{v_0^2sin2\theta}{g},"

here, "v_0" is the initial velocity of the projectile, "\theta" is the launch angle and "g" is the acceleration due to gravity.